OpenStudy (anonymous):

What Numer is between 0.5 and 5/8?

OpenStudy (sloppycanada):

5/8 = .625

OpenStudy (sloppycanada):

So a number between .5 and .625 correct?

OpenStudy (anonymous):

I belive so. Can you explain how it works?

OpenStudy (sloppycanada):

Just find the decimal of 5/8. To do this, you divide the numerator by the denominator

OpenStudy (sloppycanada):

So what's the decimal of 2/3?

OpenStudy (anonymous):

0.6?

OpenStudy (hwyl):

do you want to use fraction? \(1/2 = 0.5 = 4/8\) so \(\large \frac{4}{8}<x<\frac{5}{8} \)

OpenStudy (sloppycanada):

In 2/3, you want to expand it, so it'd be .66

OpenStudy (anonymous):

Yes.

OpenStudy (hwyl):

pay attention to what I posted

OpenStudy (sloppycanada):

Assuming all you have to do is find an integer between the two, start dividing fractions.

OpenStudy (anonymous):

Okay I just found the answers on the back of the book. Should I post them so we can figure out which one it is? All of the answers are fractions.

OpenStudy (sloppycanada):

Sure?

OpenStudy (anonymous):

a.1/4, b.8/28, c.9/16, d.19/28

OpenStudy (hwyl):